OpenStudy (anonymous):

how would you change this degrees into radians? 120° I really need to know the exact steps

OpenStudy (campbell_st):

well an easy way is to use the fact \[\pi = 180\] so 1 radian = 180/pi or 1 degree = pi/180

OpenStudy (campbell_st):

there are some basics 90 = pi/2 45 = pi/4 60 = pi/3 30 = pi/ 6 etc

OpenStudy (campbell_st):

so 120 degrees = pi/180 * 120 = (2pi)/3

OpenStudy (anonymous):